the average time it takes for a college student to complete a certain examination is 46.2 minutes.the standard deviation is 8 minutes.assume the variable is normally distributed. find the probability that a randomly selected college student will complete the examination in less than 43 minutes

1
Expert's answer
2021-11-08T20:07:33-0500

"P(X<x)=P(Z<\frac{x-\mu}{\sigma})"


"P(X<43)=P(Z<\frac{43-46.2}{8})=P(Z<-0.4)"


From table of cumulative standard normal distribution, it is found that the probability of a randomly selected college student will complete the examination in less than 43 minutes is:


"P(X<43) = P(Z<-0.4)=0.3446"